Running a small pen shop in Vancouver you purchase 100 boxes of blue pens from Ink
Inc. at a trade discount of 20%. You receive an invoice for your order on March 15 in the
amount of $2050 with terms of 2/10, n/30. You completely pay the invoice on March 25.
Your accountant has suggests that you should incorporate unit overhead costs of 30% of
cost to all of your stock. Also, you should set the price so that operating profit is 20% of
cost, if the stock is sold at a 25% discount. What is the regular selling price of a box of
fancy pens?
1
Expert's answer
2020-06-01T20:03:11-0400

1.payment on March 25 with a 2% discount with terms of 2/10, n/30:

"2050-2050\times0.02=2050-41=2009"

2.include overhead:

"2009+2009\times0.3=2009+602.7=2611.7"

3.lay profit is 20% of cost, if the stock is sold at a 25% discount:

"2611.7+(0.45-0.25)\times 2611.7=3134.04"

3.determine the price

"\frac{3134.04}{100}=31.34"